consumption=50+0.6yd
yd=y-TA+TR
TR=100
TA=ty=0.2y
I=50
G=75
NX=100
find AD equation
find consumptio
find aggregate demand
find TA
find slope of AD
draw AD curve by assuming different levels of y
1
Expert's answer
2019-10-23T09:04:42-0400

Y=C+I+G+NX

C=50+0.6(Y-TA+TR) = 50+0.6(Y-0.2Y+100)=50+0.6Y-0.12Y+60=110+0.48Y

Y = 110+0.48Y+50+75+100

Y = 335+0.48Y

0.52Y=335

Y=644.23

C = 110+0.48*644.23

C = 419.23

TA = 0.2*644.23

TA = 128.84

the slope is negative